OpenStudy (anonymous):

Suppose the area of a circle is 16pi. What is its diameter?

terenzreignz (terenzreignz):

Okay, real quick... what's the formula for the area of a circle?

OpenStudy (anonymous):

Area=pi*radius^2

terenzreignz (terenzreignz):

\[\Large \pi r ^2\] and this is equal to \(\large16\pi\), which IS the area of the given circle. Given that, can you solve for the radius r? ^_^

OpenStudy (anonymous):

8

terenzreignz (terenzreignz):

Stuck? Start with \[\large 16\pi = \pi r^2\] And divide both sides by \(\pi\)

OpenStudy (anonymous):

Thanks. :D

terenzreignz (terenzreignz):

No problem. :)

OpenStudy (anonymous):

yea and the answers 8